1.) An aqueous solution of sodium hydroxide reacts with an aqueous solution of hydrochloric acid, yeilding water. What is the balanced net ionic equation?

2.) An aqueous solution of sodium hydroxide reacts with an aqueous solution of ammonium chloride, yielding aqueous ammonia, NH3, and water. What is the balances net ionic equation?

3.) An aqueous solution of hydrochloric acid reacts with aqueous ammonia, NH3, yielding aqueous ammonium chloride. What is the balanced net ionic equation?
1
Expert's answer
2017-02-10T06:22:22-0500

1) NaOH + HCl = NaCl+H2O
2) NaOH + NH4Cl = NH3 + H2O + NaCl
3) HCl + NH3 · H2O = NH4Cl + H2O